find the total cost of a pair of jeans that costs $32 and has an 8% sales tax.
Answer by stanbon(75887)

find the total cost of a pair of jeans that costs $32 and has an 8% sales tax.
---
cost = 32 + 0.08*32 = 32(1.08) = $34.56
----------------------
Cheers,
Stan H.